Let u=<-2,1,-1, V =< - 3,2 - 1> and w=<1,3,5>. Compute :

a.) u x w,

b.) u x (w x v) and (u x w) x v.


1
Expert's answer
2021-06-25T07:23:00-0400

Let "u=<-2,1,-1>, v=<-3,2,-1>, w=<1,3,5>".

Compute:

a) "u\times w=
 \begin{vmatrix}
 i &j &k \\
 -2&1&-1\\
1&3&5
\end{vmatrix}=i (1·5 - (-1)·3) - j ((-2)·5 - (-1)·1) + k ((-2)·3 - 1·1) 

 = i (5 + 3) - j (-10 + 1) + k (-6 - 1) = <8; 9; -7>"

b)

"u\times (w\times v)=\begin{vmatrix}
 i&j&k \\
 1&3&5\\
-3&2&-1
\end{vmatrix}=i (3·(-1) - 5·2) - j (1·(-1) - 5·(-3)) + k (1·2 - 3·(-3)) 

 = i (-3 - 10) - j (-1 + 15) + k (2 + 9) ="

"=\begin{vmatrix}
 i&j&k \\
 -2&1&-1\\
-13&-14&11
\end{vmatrix}= i (1·11 - (-1)·(-14)) - j ((-2)·11 - (-1)·(-13)) + k ((-2)·(-14) - 1·(-13)) 

 = i (11 - 14) - j (-22 - 13) + k (28 + 13)= <-3; 35; 41>"

"(u×w)×v=\begin{vmatrix}
 i&j&k \\
 -2&1&-1\\
1&3&5
\end{vmatrix}=i (1·5 - (-1)·3) - j ((-2)·5 - (-1)·1) + k ((-2)·3 - 1·1) 

 = i (5 + 3) - j (-10 + 1) + k (-6 - 1) =
\begin{vmatrix}
 i&j&k\\
 8&9&-7\\
-3&2&-1
\end{vmatrix}=i (9·(-1) - (-7)·2) - j (8·(-1) - (-7)·(-3)) + k (8·2 - 9·(-3)) 

 = i (-9 + 14) - j (-8 - 21) + k (16 + 27) = <5; 29; 43>."